OpenStudy (anonymous):

-1/12 - 3/4 = A. -1/3 B. -5/6 C. 2/3 D. -4/16

OpenStudy (phi):

can you find the common denominator?

OpenStudy (anonymous):

4

OpenStudy (phi):

for 12 list out its multiples: 12,24,36 for 4: 4,8,12,16,20 look for a number on both lists

OpenStudy (anonymous):

As phi said, both 4 and 12 have a common multiple of 12. 4 multiplied by 3 equals 12, so multiply the numerator(3) by 3 as well. \[\frac{ -1 }{ 12 } *1 = \frac{ -1 }{ 12 }\]\[\frac{ 3 }{ 4 } *3 = \frac{ 9 }{ 12 }\] Now find the difference of the numerators: \[\frac{ -1 }{ 12 } - \frac{ 9 }{ 12 } = \frac{ (-1) - 9 }{ 12}\]